[Hallc_running] [New Logentry] RC Report -- 09/06/22

dhamilto at jlab.org dhamilto at jlab.org
Tue Sep 6 08:50:02 EDT 2022

Logentry Text:
We are just in the process of completing production running at 2.5 uA on LD2 for the final kinematic setting in the run plan (Q2=1.60, W=3.08 Th_shms=6.00). There have been no major issues: beam was down for a couple of hours during Swing due to a quadrupole mismatch and we had an HMS Q1 trip during owl (it came back after a reset).

It looks like beam may be available earlier than expected tomorrow after the maintenance day down. We should expect to hear more later today. Ops have requested that we finish at 8.30am rather than 8am on Friday.

The modified run plan involves:

1) Switching to the Q2=3.85, W=2.62, Th_SHMS=8.00 setting to take data at 12 uA on LD2 until the maintenance day down (we also need around 200 mC of dummy target data).
2) Low current BCM calibration using the Faraday cup when beam is restored on Wednesday (coordinate with Dave Mack).
3) Changing to the the Q2=3.85, W=2.62, Th_SHMS=10.00 setting to take data at 38 uA on LD2 for around two shifts (we also need around 60 mC of dummy target data at this setting).
4) One last change back to the Q2=3.85, W=2.62, Th_SHMS=8.00 setting to take data at 12 uA on LD2 until Friday morning.

Let's have a short meeting today at 11am to discuss.
Meeting ID: 160 710 0885 - Passcode: 595071


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4038313
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220906/096b1656/attachment.html>

More information about the Hallc_running mailing list